OpenStudy (anonymous):

What is the value of x in the equation? (-17) + 20 – (9 × 8) ÷ 3 = x

OpenStudy (anonymous):

i got -27???

OpenStudy (anonymous):

(-17) + 20 – (72) ÷ 3 = x ( parenthesis first) (-17) + 20 –(24) = x ( following the MDAS rule we should multiply of divide first whichever comes first on the left :) then you can now simplify it by adding or subtracting whichever comes first on the left :) 3 -24 = x x= -21

OpenStudy (anonymous):

@mowrey and @JamenS did you followed the rule of PEMDAS ? P arenthesis E xponent M ultiplication D ivision A ddition and S ubtarction ??

OpenStudy (anonymous):

oh i see...thx

OpenStudy (anonymous):

welcome :)